### v2

# Positioning is All You Need

Abstract
One can drive safely with a GPS without memorizing a world map (not to mention the dark regions that humans have never explored). Such a locality-based attention mechanism has a profound implication on our understanding of how the brain works. This paper refines the existing embodied cognition framework by turning the locality from a constraint to an advantage. Analogous to GPS-based navigation, positioning represents a computationally more efficient solution to flexible behaviors than reconstruction. This simple intuition implies that positioning is all you need to understand cognitive functions generated by hippocampal-neocortical systems. That is, the neocortex generates thousands of local maps whose indexing is maintained by the hippocampus. Geometrically, we present a simple manifold positioning framework to explain the principle of localized embodied cognition. The positioning operation implemented by the attention mechanism can be interpreted as a nonlinear projection linking the discovery of local subspace structure by the neocortex (a sensorimotor machine interacting with the world locally) for the navigation task in mind $without$ discovering global manifold topology.
Introduction
Sensory organs attached to the mammalian brain can access only a localized region in space at any specific time. Such spatio-temporal locality of sensory observation has a profound implication for the mental construction of the internal world (a.k.a., internal representation  (?)) including consciousness by the mammalian brain. Nature has discovered a clever solution to learn a global representation of the physical world with a local receptive field of sensory cortex by flexible behavior, including spatial navigation. How do mammalian brains encode the spatial information about the physical world into the population of spike sequences in the temporal domain? Such a question is at the holy grail of neuroscience whose importance to cognitive science is equivalent to relativity theory in modern physics.
To formulate the above problem geometrically, we note that the observation data acquired by localized sensory organs can be abstracted by the projection of a manifold embedded in a high-dimensional to a low-dimensional subspace. It is well recognized that nonlinear manifold, despite the sophisticated topology globally, is locally isomorphic to a Euclidean space. Such fundamental property has been the inspiration for the class of nonlinear dimensionality reduction techniques (e.g., IsoMAP  (?) and locally linear embedding  (?)). An important new insight brought by this work is that we do not need to reconstruct the global manifold but only need to know the $position$ within the global coordinates for goal-oriented behavior. This insight is in alignment with intelligence without representation  (?) and animate vision  (?) which attempt to put perception and action on an equal footing. Using GPS as the metaphor, we only need access to a local map without the need to memorize the global map anywhere and anytime. Such locality principle, when combined with embodied cognition, is a fundamental advantage that has been exploited by nature during the evolution of mammalian brains.
Through co-evolution of the hippocampus and the neocortex, nature solves the manifold positioning problem by the simple attention mechanism  (?), which can be mathematically modeled by context  (?) and computationally implemented by the transformer architecture  (?). At the level of individual neurons, attention or selection of firing patterns is implemented by the nonlinear effects of inhibitory networks  (?). At the level of cortical columns, canonical circuits implementing predictive processing  (?) are responsible for context-dependent modulation of prediction errors. At the level of cortical areas, hippocampal indexing theory  (?, ?) appears to be the most comprehensive treatment of modeling hippocampal-neocortical interactions  (?). By serving as the neocortex’s librarian (pp. 285,  (?), the hippocampus uses an indexing-based positioning strategy to dynamically stabilize the sensorimotor activities generated by the neocortex.
In this paper, we provide a geometric interpretation of the positioning/indexing strategy underlying hippocampal-neocortical systems. It is shown that the problem of navigating on an unknown manifold in an arbitrary dimension can be solved by “think globally, fit locally”  (?) - i.e., to exploit the local subspace constraint of a nonlinear manifold, the only solution that is universal and generalizable is to explicitly learn a nonlinear mapping that connects the local geometry with the global topology. Since we are not interested in manifold reconstruction (e.g., the dark areas of the world map), an indexing or positioning strategy is sufficient for “navigating” from one place to another. This intuition is consistent with “using the world as its own model”  (?) but we are taking one step forward by noting that we only need a localized version of the world model due to the locality constraint of our sensory organs. Using the map analogy from a thousand brain theory  (?), we need a librarian (i.e., the hippocampus) to maintain efficient indexing and retrieval of 150,000 maps generated by the neocortex. This librarian can easily manage more maps as the intelligence evolves - note that the neocortex has a virtually infinite capacity thanks to its small-world network organization  (?); while the network organization of the hippocampus does not change much from mammals to humans.
Toy Example
To facilitate the illustration of our core idea, we introduce a toy example based on marrying place cells as discovered in the hippocampus  (?) with the two-spiral dataset as widely used in manifold learning literature  (?), as shown in Fig. 1. Even though the maze experiments used in studying rodent models have not used such a spiral setting  (?), we think it is plausible to assume that mammalian brains such as rodents can easily manage a maze with a spiral shape. The deeper implication of this toy example is that if place cells are interpreted as spatial context, we can generalize this hypothesized experiment to higher dimensions. For example, if object recognition is formulated as a maze problem in a higher-dimensional space, how can we despositionign an intelligence machine for navigation?
a)                                                      b)
Figure 1: The place cells discovered in the hippocampus provide spatial context for the task of navigation. Such contextual modeling can be exploited to navigate the local Euclidean space of an arbitrary manifold by “memorizing” its position in the global coordinates.
In conventional wisdom such as IsoMAP  (?) and locally linear embedding (LLE)  (?), the objective is to exploit the local subspace constraint of the manifold by nonlinear dimensionality reduction. The drawback of this class of solutions is that they do not generalize well to arbitrary manifolds, especially in higher dimensions. The fundamental reason behind this failure can be explained by intelligence without representation  (?) - under the framework of embodied cognition  (?), perception and action have an equal footing. From the evolutionary perspective, goal-oriented behavior is the foundation of natural intelligence. Therefore, $navigation$ on the manifold is easier than $reconstruction$ of the manifold; and nature prefers simple but not simpler solutions. A moment of thought can tell that a place cell-based navigation strategy can be easily generalized to a manifold of arbitrary dimension or topology. The job of indexing or map-keeping will become more involved but it is still manageable by a librarian.
Computational Modeling of Hippocampal-Neocortex Systems
We advocate a new conceptual framework called “localized embodied cognition” with the locality principle in mind from problem formulation to solution algorithms. We still use the world “as its own model”, but the way local embodied cognition never attempts to solve a global problem locally. An organism needs to couple sensory with motor systems to achieve various goals but never stores knowledge about the global environment. Instead, the bodily interactions with the environment (situatedness) are persistently maintained by the neocortex (a single book)’s interaction with the hippocampus (a large library). The cognitive map in the hippocampus offers a situational context to interpret the sensorimotor interaction occurring in the neocortex. In this section, we first present computational models for characterizing indexing operation in the hippocampus and sensorimotor interaction in the neocortex and then connect them with a universal attention mechanism.
a)                                                      b)
Figure 2: Computational modeling of hippocampal-neocortical coupling.
Hippocampal Indexing Theory.
The key idea underlying hippocampal indexing theory  (?, ?) is that the hippocampus is “functional designed and anatomically situated” to handle the diverse neocortical activities generated by sensorimotor interactions. By projecting back to the neocortex, the hippocampus stores a collection of indexes that can serve as the context for the storage and retrieval of episodic memory  (?). Recent findings indicate a more complex interaction between hippocampal functioning and episodic memory, with an emphasis on the cognitive map interpretation  (?). This newer perspective considers the hippocampus as creating spatial maps (cognitive maps  (?)) for memory formation. The theory also recognizes the hippocampus’s role in binding different elements of an experience, including time and space, into a cohesive memory  (?). Furthermore, the relationship between place cells in the hippocampus and memory formation suggests that these cells may function both as part of a cognitive map and as index cells for memory  (?).
Neocortex as a Sensorimotor Machine.
One of the biggest mysteries in studying cortical columns is that they are structures without a function  (?). Instead, several canonical microcircuits such as Douglas-Martin model  (?) and Haeusler-Maass model  (?) have been studied for predictive coding in the literature. These models have revealed a remarkable correspondence between the microcircuitry of the cortical column and the connectivity required by predictive coding. In Douglas-Martin model  (?), information flow through the cortical column is characterized by a stereotypical pattern of fast excitation followed by slower and longer-lasting inhibition in the cat visual system. The three neuronal populations receive thalamic drive and amplify transit thalamic inputs to generate sustained activities while maintaining a balance between excitation and inhibition. This model was further extended by Haeusler and Maass in  (?), which closely resembles the canonical circuits required by predictive coding. The feedforward prediction errors from a lower cortical level arrive at granular layers and are passed to excitatory and inhibitory interneurons in supragranular layers, encoding expectations. Meanwhile, the connections between excitatory and inhibitory neurons in supragranular layers enable deep pyramidal cells and excitatory interneurons to generate context-dependent feedback predictions, which descend to a lower hierarchical level.
Hippcampal-Neocortical Interaction
. The coupling of the hippocampus with the neocortex is based on the principle of localized embodied cognition (please refer to Fig. 2). Following the spirit of intelligence without representation  (?), we use the localized world as its own model which nicely fits the functional design of the neocortex (i.e., detecting regularity in stimuli by sensorimotor interaction). To accommodate the constantly changing properties of the global world model, an organism needs a cognitive map for navigation. This map can be mathematically abstracted by a nonlinear mapping between local subspace geometry as discovered by the neocortex and global topology as maintained by the hippocampus. Instead of reconstructing the entire manifold, we argue it is sufficient to retrieve its local projection based on the context information. This nonlinear mapping is essentially an attention mechanism that can be implemented by a sparse and distributed memory (SDM)  (?) (attention approximates SDM  (?)) or an associative memory (e.g., modern Hopfield network  (?) is intrinsically connected with the transformer architecture  (?)). SDM-based indexing/positioning supports continual learning  (?). Real-world map-based navigation serves as a perfect example of the plausibility of this manifold positioning strategy. Why can we safely drive around using GPS without memorizing the entire world? Because at any time and location, we only need access to a localized version of the world map for navigation purposes. Note that there is no need to reconstruct a map for an area if no one ever goes there, which is consistent with goal-oriented behavior in natural intelligence.
Applications into Cognitive Science
In this section, we revisit four widely studied cognitive tasks from the new perspective of locally navigating on a globally unknown manifold. The unifying theme is that positioning is all you need to understand how the human brain solves these problems by a universal cortical processing algorithm, as envisaged by Vernon Mountcastle in 1957  (?).
Spatial Navigation.
Cognitive maps organize knowledge to support flexible and adaptive behavior. Among various behavioral tasks, spatial navigation is among the most well-studied in the literature  (?, ?). How do mammalian brains encode 2D physical space information into 1D spike sequence? This question is at the core of understanding the relationship between space and time in its simplest form. The representation of spatial information (Euclidean or non-Euclidean) plays an essential role in the building of cognitive maps. A graph-based representation has been widely adopted for encoding the physical space into action-related state space. A fundamental problem is the aliasing problem because identical sensory observations can occur in different spatial locations  (?). The aliasing problem is solved by attention in the hippocampus where the encoding of spatio-temporal context (index to the local map or position on the global map) is handled by CA3-CA1 collateral system (pp. 326  (?)). The consolidation of semantic memory (analogy to global manifold topology) from episodic memory (analogy to local Euclidean geometry) eliminates the spatio-temporal context to support context-dependent inference where context is defined by the attention window associated with different spatial and temporal locations.
Motor Control.
Motor control refers to the process of physically interacting with objects in the world and manipulating them toward completing a specific task/goal (e.g., reach and grasp). In motor control and trajectory planning  (?), internal models consist of: 1) a forward component responsible for predicting sensory consequences of the executed motor commands; and 2) an inverse dynamics model that calculates forward motor commands from the desired trajectory information. Both components require an attention-guided and action-oriented body map  (?) that is centered on the controlled object. The generalization property of motor controls lies in the transformation from one coordinate to another. For example, the task of grasping a water bottle differs from that of grasping a coffee cup at the surface; but the principle of motor commands is the same except for the change of coordinates (from egocentric for spatial navigation to allocentric for motor control). Using the library as the metaphor again, what the hippocampus needs to adapt is to $re-position$ the neocortex - i.e., from one book of grasping without a handle to another with a handle. Unlike navigation, object manipulation does not require path integration for trajectory planning because the scale of motor control is the same as that of the reference frame centered on the object.
Object Detection.
In the spatial navigation of rodents, the environment or place learning is achieved by ego-motion of animals, which dynamically changes the bearing and sketch maps  (?). Such behavior becomes more sophisticated as the evolution of the visual cortex helps locate salient objects (a.k.a. landmarks) in the environment  (?). From a manifold positioning perspective, we argue that there is an appealing analogy between the integrated map for spatial navigation and the shape description for animate vision  (?). That is, the problem of object/landmark detection can also be solved by the inferior temporal (IT) cortex  (?) assuming that the spatial-temporal context is provided by the hippocampus. Mathematically, a place cell can be abstracted by a Dirac function $\delta(x,y)$ that only fires when the organism is in the neighborhood of $(x,y)$. By analogy, a gnostic cell (a.k.a. concept cell  (?, ?) can fire if and only if an object is detected by the neocortex in a latent space (e.g., generalized Hough transform  (?)). The spatial attention mechanism implemented by the gaze becomes critically important here because it reduces the computational burden by indexical reference  (?). More importantly, positioning of the gaze direction is naturally required by the visual cortex to recognize an object based on the detection result  (?).
Language Comprehension.
Action-oriented representation facilitates understanding the constitution between language and action in the theory of embodied language  (?). By restoring the central role played by the motor system in language processing, we can gain deeper insight into the essence of language from a motor resonance or action simulation perspective  (?). The class of referential motor resonance is the source of abstract thoughts or covert movements in our mental world, which are hierarchically organized. The relationship between hierarchical motor control  (?) and the nested structure of natural languages  (?) can be better understood from a hierarchical extension of the attention perspective. Both language and motor control demonstrate compositionality, whereby complex structures are built from simpler elements according to combinatorial rules. In natural language, sentences are composed of phrases composed of smaller units such as words and morphemes. This syntactic structure manifests as the ability to generate an infinite variety of sentences from a finite set of words and grammatical rules. Similarly, in motor control, movements can be decomposed into smaller units, such as motor primitives, which are organized hierarchically to generate complex behaviors. Moreover, movements can comprise sequences of simpler actions or motor primitives, combined to produce more complex behaviors. The latest study on relating neural representation of the hippocampal formation (e.g., Tolman-Eichenbaum Machine  (?)) to transformer  (?) seems to support that attention is needed for both spatial navigation and language comprehension. The curios position formula of tokens in the transformer architecture using sine and cosine functions seems the de-aliasing result that is conceptually similar to the omnidirectional place cells in cognitive maps (pp. 330,  (?)).